OpenStudy (anonymous):

The members of your high school band want to purchase new jackets. The silk-screen shop charges a $23.2 set up fee and $29.2 per jacket. If you order more than 45.6 jackets, the price per jacket is reduced to $24.2. Write a compound function that gives the cost C for an order of of x jackets and sketch the graph of the function.

OpenStudy (danjs):

X = number of jackets if x>45.6 then Cost = 23.2 + 24.2x if x< or =45.6 then Cost = 23.2 + 29.2

OpenStudy (danjs):

So if you buy MORE than 45.6 jackets, you are charged 23.2 set up fee plus 24.2 per jacket X

OpenStudy (danjs):

If you buy LESS then 45.6 or exactly 45.6 jackets, the cost is still the 23.2 set up fee, plus 29.2 per jacket this time

OpenStudy (anonymous):

okay so

OpenStudy (danjs):

Cost = Y Jackets = X Y = 23.2 + 29.2*X when X< or = 45.6 Y = 23.2 + 24.2*X when X>45.6

OpenStudy (anonymous):

C =29.2x + 23.2, x ≤ 45.6 and 24.2x + 23.2, x > 45.6

OpenStudy (anonymous):

C =29.2x + 23.2, x ≤ 45.6 and 24.2x + 23.2, x > 45.6

OpenStudy (anonymous):

why is it doing that?

OpenStudy (danjs):

reload page

OpenStudy (danjs):

copy and paste special characters is messed up

OpenStudy (danjs):

the first one is it

OpenStudy (anonymous):

the first one I posted

OpenStudy (danjs):

yeah both are the same, looks good

OpenStudy (anonymous):

do you have time to help me with more questions?

OpenStudy (danjs):

i can do one or 2 more, then have to go for a little bit

OpenStudy (anonymous):

Use the table x 1 2 3 4 5 6 f(x) 1 6 5 2 1 3 g(x) 6 2 5 2 2 4 to evaluate (g ◦ f)(3).

OpenStudy (anonymous):

s Use the table to evaluate (f ◦ g)(6). Use the table to evaluate (f ◦ f)(1).

OpenStudy (anonymous):

@DanJS